Find the present values of the following cash flow streams at an 8% discount rate

UNEVEN CASH FLOW STREAM


Find the present values of the following cash flow streams at an 8% discount rate. Round your answers to the nearest cent.


012345

Stream A $  
Stream A$0$150$450$450$450$300
Stream B$0$300$450$450$450$150

Stream B $  


What are the PVs of the streams at a 0% discount rate?

Stream A $  

Stream B $  
Answer #1

Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

At 8%:

A=150/1.08+450/1.08^2+450/1.08^3+450/1.08^4+300/1.08^5

=$1416.85(Approx)

B:=300/1.08+450/1.08^2+450/1.08^3+450/1.08^4+150/1.08^5

=$1453.66(Approx)

At 0%:

A=150+450+450+450+300=$1800

B=300+450+450+450+150=$1800

[Please note that intermediate calculations have not been rounded off].